OpenStudy (anonymous):

giving medal!!!!!!!!help?? In 1990, the population of Delray Beach, Florida was 47,184. In 2000, the population was 60,020. What is the percent of change in population? Round your answer to the nearest tenth, if necessary. a. 21.4% c. –27.2% b. 27.2%` d. 78.6%

OpenStudy (anonymous):

give it then ill answer

OpenStudy (anonymous):

First, you'll want to find the change in population. This is equal to the final population minus the initial population: 60,020-47,184 Once you find this, then divide it by the original population, 47,184, and multiply by 100%: \[\frac{ 60020-47184 }{ 47184 } \times 100 percent\]

OpenStudy (anonymous):

From 47 lb to 52 lb: let say big number is 100% and you have 52 is 100% 47 is ? % ?% = 47 x 100 / 52 = 90.3846.....% now simply minus 100% - 90.3846....% = 9.6% ( letter A ) this a way how to do it